# Hyperelliptic four-manifolds defined by vector-colorings of simple polytopes

### v2

@@ 3. A criterion when $N(P,\Lambda)\simeq S^{4}$
Definition 3.2 (see [V71, D83, D08]).
Define a space
@@
Any vector-coloring $\Lambda$ defines an epimorphism $\varphi_{\Lambda}\colon\mathcal{G}(P)\to\mathbb{Z}_{2}^{r}$.
@@
If $\pi_{1}(N(P,\Lambda))=0$, then ${\rm Ker\,}\varphi$} is generated by kernels of the mappings $\langle g\rho_{i_{1}}g^{-1},\dots g\rho_{i_{n}}g^{-1}\rangle\simeq\mathbb{Z}_{2}^{n}\to\mathbb{Z}_{2}^{r}$, $g\rho_{i_{s}}g^{-1}\to\Lambda_{i_{s}}$, for all elements $g\in\mathcal{G}(P)$ and all vertices $F_{i_{1}}\cap\dots\cap F_{i_{n}}$ of $P$.
Consider the composition $\mathcal{G}(P)\to\mathbb{Z}_{2}^{m}\to\mathbb{Z}_{2}^{r}$, where the first mapping $\pi\colon\rho_{i}\to\boldsymbol{e}_{i}$ is the abelenization homomorphism, and the second mapping $\Lambda$ is a linear mapping defined by the condition $\boldsymbol{e}_{i}\to\Lambda_{i}$. We have $\pi^{-1}({\rm Ker\,}\Lambda)={\rm Ker\,}\varphi$ and
$\pi\langle g\rho_{i_{1}}g^{-1},\dots g\rho_{i_{n}}g^{-1}\rangle=\pi\langle\rho_{i_{1}},\dots\rho_{i_{n}}\rangle=\langle\boldsymbol{e}_{i_{1}}\dots\boldsymbol{e}_{i_{n}}\rangle.$
In particular, the restrictions of $\pi$ to $\langle g\rho_{i_{1}}g^{-1},\dots g\rho_{i_{n}}g^{-1}\rangle$ is injective for any vertex $v$ and any $g\in\mathcal{G}(P)$. Also
$\pi^{-1}({\rm Ker}\,\Lambda\left.\right|_{\langle\boldsymbol{e}_{i_{1}}\dots\boldsymbol{e}_{i_{n}}\rangle})={\rm Ker}\,\varphi\left.\right|_{\langle g\rho_{i_{1}}g^{-1},\dots g\rho_{i_{n}}g^{-1}\rangle}.$
Since ${\rm Ker\,}\varphi$} is generated by kernels ${\rm Ker}\,\varphi\left.\right|_{\langle g\rho_{i_{1}}g^{-1},\dots g\rho_{i_{n}}g^{-1}\rangle}$, the group $\pi({{\rm Ker\,}\varphi})={\rm Ker}\,\Lambda$ is generated by the subgroups $\pi({\rm Ker}\,\varphi\left.\right|_{\langle g\rho_{i_{1}}g^{-1},\dots g\rho_{i_{n}}g^{-1}\rangle})={\rm Ker}\,\Lambda\left.\right|_{\langle\boldsymbol{e}_{i_{1}}\dots\boldsymbol{e}_{i_{n}}\rangle}$ for all vertices $F_{i_{1}}\cap\dots\cap F_{i_{n}}$ of $P$, that is
(1) ${\rm Ker}\,\Lambda=\sum\limits_{v\in P}{\rm Ker}\,\Lambda\left.\right|_{\langle\boldsymbol{e}_{i_{1}}\dots\boldsymbol{e}_{i_{n}}\rangle}$
If $N(P,\Lambda)$ is a closed topological manifold, then by [E24, Theorem 5.1] at each vertex $F_{i_{1}}\cap\dots\cap F_{i_{n}}$ different vectors among $\{\Lambda_{i_{1}},\dots,\Lambda_{i_{n}}\}$ are linearly independent. In particular, ${\rm Ker}\,\Lambda\left.\right|_{\langle\boldsymbol{e}_{i_{1}}\dots\boldsymbol{e}_{i_{n}}\rangle}$ is generated by the vectors $\boldsymbol{e}_{i_{a}}+\boldsymbol{e}_{i_{b}}$ with $\Lambda_{i_{a}}=\Lambda_{i_{b}}$. We have $m]=\{1,\dots,m\}=S_{1}\sqcup\dots\sqcup S_{t}$, where $\Lambda_{i}=\Lambda_{j}$ (equivalently, $\boldsymbol{e}_{i}+\boldsymbol{e}_{j}\in{\rm Ker}\,\Lambda$) if and only if $i$ and $j$ lie in the same set $S_{k}$. Assume that there is a linear dependence $\Lambda_{j_{1}}+\dots+\Lambda_{j_{l}}=0$, where all the vectors $\{\Lambda_{j_{1}},\dots,\Lambda_{j_{l}}\}$ are pairwise different, that is $e_{j_{p}}\in S_{i_{p}}$, where $i_{p}\neq i_{q}$ if $j_{p}\neq j_{q}$. Then $\boldsymbol{e}_{j_{1}}+\dots+\boldsymbol{e}_{j_{l}}\in{\rm Ker}\,\Lambda$, and by (1)
@@
Indeed, $G_{1}\cap\dots\cap G_{M-1}$ is an (nM+1)(n-M+1)-RHDRHD. In particular, it is nonempty. But the intersection of any $n+1$ facets of $P$ is empty, hence $M-1\leqslant n$. Each $k$-face of $C(P,\Lambda)$ is a connected component of $G_{i_{1}}\cap\dots\cap G_{i_{n-k}}$, hence by Lemma 3.10 it is a $k$-RHDRHD if nk<Mn-k<M, and a $k$-RHDRHD if nk=Mn-k=M. ∎
@@
For $n=2$ the statement is trivial.
@@
Let $n=3$. If $M=1$, then $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,1)$. If $M=2$, then $G_{1}$ and $G_{2}$ are $2$-disks such that $\partial G_{1}=\partial G_{2}=G_{1}\cap G_{2}$ is a circle. Hence, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,2)$. If $M=3$, then $G_{1}$, $G_{2}$ and $G_{3}$ are $2$-disks such that $G_{i}\cap G_{j}$ is a simple path for any $i\neq j$, and all these three paths have common ends, which form the 0-sphere $G_{1}\cap G_{2}\cap G_{3}$. Thus, the $1$-skeleton of $\mathcal{C}(P,\Lambda)$ is a theta-graph, and $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,3)$. If $M=4$, then each $G_{i}$ is a disk, the intersection $G_{i}\cap G_{j}$ of each two such disks is a simple path, the intersection of any three such disks is a point, and the intersection of all the $4$-disks is empty. Then the $1$-skeleton of $\mathcal{C}(P,\Lambda)$ is a $K_{4}$-graph, and $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,3)$.
@@
Let $n=4$. If $M=1$, then $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,1)$. If $M=2$, then $G_{1}$ and $G_{2}$ are $3$-disks such that $\partial G_{1}=\partial G_{2}=G_{1}\cap G_{2}$ is a $2$-sphere. Hence, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,2)$. If $M=3$, then $G_{1}$, $G_{2}$ and $G_{3}$ are $3$-disks, such that $G_{i}\cap G_{j}$ is a $2$-disk for $i<j$, and $G_{1}\cap G_{2}\cap G_{3}$ is a circle. We have $G_{1}\cup G_{2}$ is a $3$-disk with a circle $G_{1}\cap G_{2}\cap G_{3}=\partial(G_{1}\cap G_{2})$ on its boundary, and the complement to this disk is the interior of $G_{3}$. Hence, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,3)$. If $M=4$, then $G_{1}$, $G_{2}$, $G_{3}$, $G_{4}$ are $4$-disks, $G_{i}\cap G_{j}$ is a $2$-disk for any $i<j$, $G_{i}\cap G_{j}\cap G_{k}$ is a simple path for any $i<j<k$, and all these four paths have the common ends, which form the 0-sphere $G_{1}\cap G_{2}\cap G_{3}\cap G_{4}$. Then $G_{1}\cup G_{2}\cup G_{3}$ is a $3$-disk with the theta-graph on the boundary formed by paths $G_{i}\cap G_{j}\cap G_{4}$, $i<j<4$. There is the path $G_{1}\cap G_{2}\cap G_{3}$ inside this disk connecting the vertices of the theta-graph and each $G_{i}\cap G_{j}$ is a $2$-disk bounded by the paths $G_{1}\cap G_{2}\cap G_{3}$ and $G_{i}\cap G_{j}\cap G_{4}$. The complement to $G_{1}\cup G_{2}\cup G_{3}$ in $\partial P$ is the interior of $G_{4}$. Thus, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,4)$. If $M=5$, then each $G_{i}$ is a $3$-disk, each intersection $G_{i}\cap G_{j}$, $i<j$, is a $2$-disk, each intersection $G_{i}\cap G_{j}\cap G_{k}$, $i<j<k$, is a simple path, each intersection $G_{i}\cap G_{j}\cap G_{k}\cap G_{l}$, $i<j<k<l$, is a point, and $G_{1}\cap G_{2}\cap G_{3}\cap G_{4}\cap G_{5}=\varnothing$. Then $G_{1}\cup G_{2}\cup G_{3}\cup G_{4}$ is a $3$-disk with a $K_{4}$-graph on its boundary formed by vertices $G_{i}\cap G_{j}\cap G_{k}\cap G_{5}$ and edges $G_{i}\cap G_{j}\cap G_{5}$. There is a point $G_{1}\cap G_{2}\cap G_{3}\cap G_{4}$ inside this disk such that the face structure in the disk $G_{1}\cup G_{2}\cup G_{3}\cup G_{4}$ is a cone over the face structure on its boundary with this apex. Also the complement to this disk is the interior of $G_{5}$. Thus, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,5)\simeq\partial\Delta^{4}$.
@@
Let $\Lambda$ be a vector-coloring of rank $r$ of a simple polytope of dimension $n\leqslant 4$. Then $N(P,\Lambda)\simeq S^{n}$ if and only if $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,r)$.
@@ 4. Hyperelliptic manifolds $N(P,\Lambda)$
Let $\Lambda$ be a vector-coloring of rank $r$ of a simple $n$-polytope $P$, $n\leqslant 4$. Then an involution $\tau\in\mathbb{Z}_{2}^{r}$ is hyperelliptic (that is $N(P,\lambda)\simeq S^{n}$) if and only if $\tau$ is special (that is $\mathcal{C}(P,\Lambda_{\tau})\simeq\mathcal{C}(n,r-1)$).
@@ 4.1. Hamiltonian subcomplexes
We call a subcomplex $\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ Hamiltonian, if each $q$-skeleton of $\mathcal{C}(P,c_{1})$ belongs to the $(q+1)$-skeleton of $\mathcal{C}(P,c_{2})$.
@@
A subcomplex $\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ is Hamiltonian if and only if any two defining (n2)(n-2)-facets are disjoint.
@@
Let any two defining (n2)(n-2)-facets be disjoint. If $c_{2}(G_{i_{1}})=c_{2}(G_{i_{2}})=c_{2}(G_{i_{3}})$ and $i_{1}<i_{2}<i_{3}$, then either one of the intersections $G_{i_{1}}\cap G_{i_{2}}$ and $G_{i_{1}}\cap G_{i_{3}}$ is empty, or each intersection is a disjoint union of defining faces and any two faces from different unions are also disjoint. Then
@@
If $c_{2}(G_{i_{1}})=c_{2}(G_{i_{2}})\neq c_{2}(G_{i_{3}})=c_{2}(G_{i_{4}})$ and $i_{1}\neq i_{2}$, $i_{3}\neq i_{4}$, then either one of the intersection $G_{i_{1}}\cap G_{i_{2}}$ and $G_{i_{3}}\cap G_{i_{4}}$ is empty, or each intersection is a disjoint union of defining faces and any two faces from different unions are also disjoint. Then
@@
Now let the subcomplex be Hamiltonian. Consider two defining faces. If they are connected components of the same intersection $G_{i_{1}}\cap G_{i_{2}}$, then they are disjoint by definition. If they belong to different intersections $G_{i_{1}}\cap G_{i_{2}}$ and $G_{i_{3}}\cap G_{i_{4}}$, then there are two possibilities. First possibility is $c(G_{i_{1}})=c(G_{i_{2}})=c(G_{i_{3}})=c(G_{i_{4}})$. Then without loss of generality we can assume that $i_{3}\notin\{i_{1},i_{2}\}$. Hence, $G_{i_{1}}\cap G_{i_{2}}\cap G_{i_{3}}=\varnothing$ by Lemma 4.4(3) and $(G_{i_{1}}\cap G_{i_{2}})\cap(G_{i_{3}}\cap G_{i_{4}})=\varnothing$. The second possibility is $c(G_{i_{1}})=c(G_{i_{2}})\neq c(G_{i_{3}})=c(G_{i_{4}})$. Then $(G_{i_{1}}\cap G_{i_{2}})\cap(G_{i_{3}}\cap G_{i_{4}})=\varnothing$ by Lemma 4.4(3). This finishes the proof. ∎
@@
A set $\mathcal{S}=\{M_{1}^{n-2},\dots,M_{q}^{n-2}\}$ of pairwise disjoint (n2)(n-2)-faces of a complex $\mathcal{C}(P,c_{1})$ is a set of defining faces of a Hamiltonian subcomplex $\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ if and only if for each two facets $G_{i}\neq G_{j}$ of $\mathcal{C}(P,c_{1})$ admitting a sequence of facets $G_{i}=G_{i_{1}}$, $G_{i_{2}}$, $\dots$, $G_{i_{p}}=G_{j}$ such that for all $s$ the intersection $G_{i_{s}}\cap G_{i_{s+1}}$ contains a face from $\mathcal{S}$ either $G_{i}\cap G_{j}=\varnothing$ or each connected component of $G_{i}\cap G_{j}$ belongs to $\mathcal{S}$.

### v3

@@ 3. A criterion when $N(P,\Lambda)\simeq S^{4}$
Definition 3.2.
(see [V71, D83, D08])Define a space
@@
Any vector-coloring $\Lambda$ of rank $r$ defines an epimorphism $\varphi_{\Lambda}\colon\mathcal{G}(P)\to\mathbb{Z}_{2}^{r}$. The proof of the following fact is straightforward.
@@
If $\pi_{1}(N(P,\Lambda))=0$, then ${\rm Ker\,}\varphi$} is generated by kernels of the mappings $\langle g\rho_{i_{1}}g^{-1},\dots,g\rho_{i_{n}}g^{-1}\rangle\simeq\mathbb{Z}_{2}^{n}\to\mathbb{Z}_{2}^{r}$, $g\rho_{i_{s}}g^{-1}\to\Lambda_{i_{s}}$, for all elements $g\in\mathcal{G}(P)$ and all vertices $F_{i_{1}}\cap\dots\cap F_{i_{n}}$ of $P$.
Consider the composition $\mathcal{G}(P)\to\mathbb{Z}_{2}^{m}\to\mathbb{Z}_{2}^{r}$, where the first mapping $\pi\colon\rho_{i}\to\boldsymbol{e}_{i}$ is the abelianization homomorphism, and the second mapping $\Lambda$ is a linear mapping defined by the condition $\boldsymbol{e}_{i}\to\Lambda_{i}$. We have $\pi^{-1}({\rm Ker\,}\Lambda)={\rm Ker\,}\varphi$ and
$\pi\langle g\rho_{i_{1}}g^{-1},\dots,g\rho_{i_{n}}g^{-1}\rangle=\pi\langle\rho_{i_{1}},\dots,\rho_{i_{n}}\rangle=\langle\boldsymbol{e}_{i_{1}},\dots,\boldsymbol{e}_{i_{n}}\rangle.$
In particular, the restriction of $\pi$ to $\langle g\rho_{i_{1}}g^{-1},\dots,g\rho_{i_{n}}g^{-1}\rangle$ is injective for any vertex $v$ and any $g\in\mathcal{G}(P)$. Also
$\pi^{-1}({\rm Ker}\,\Lambda\left.\right|_{\langle\boldsymbol{e}_{i_{1}},\dots,\boldsymbol{e}_{i_{n}}\rangle})={\rm Ker}\,\varphi\left.\right|_{\langle g\rho_{i_{1}}g^{-1},\dots,g\rho_{i_{n}}g^{-1}\rangle}.$
Since ${\rm Ker\,}\varphi$} is generated by kernels ${\rm Ker}\,\varphi\left.\right|_{\langle g\rho_{i_{1}}g^{-1},\dots,g\rho_{i_{n}}g^{-1}\rangle}$, the group $\pi({{\rm Ker\,}\varphi})={\rm Ker}\,\Lambda$ is generated by the subgroups $\pi({\rm Ker}\,\varphi\left.\right|_{\langle g\rho_{i_{1}}g^{-1},\dots,g\rho_{i_{n}}g^{-1}\rangle})={\rm Ker}\,\Lambda\left.\right|_{\langle\boldsymbol{e}_{i_{1}},\dots,\boldsymbol{e}_{i_{n}}\rangle}$ for all vertices $F_{i_{1}}\cap\dots\cap F_{i_{n}}$ of $P$, that is
(1) ${\rm Ker}\,\Lambda=\sum\limits_{v\in P}{\rm Ker}\,\Lambda\left.\right|_{\langle\boldsymbol{e}_{i_{1}},\dots,\boldsymbol{e}_{i_{n}}\rangle}$
If $N(P,\Lambda)$ is a closed topological manifold, then by [E24, Theorem 5.1] at each vertex $F_{i_{1}}\cap\dots\cap F_{i_{n}}$ different vectors among $\{\Lambda_{i_{1}},\dots,\Lambda_{i_{n}}\}$ are linearly independent. In particular, ${\rm Ker}\,\Lambda\left.\right|_{\langle\boldsymbol{e}_{i_{1}},\dots,\boldsymbol{e}_{i_{n}}\rangle}$ is generated by the vectors $\boldsymbol{e}_{i_{a}}+\boldsymbol{e}_{i_{b}}$ with $\Lambda_{i_{a}}=\Lambda_{i_{b}}$. We have $m]=\{1,\dots,m\}=S_{1}\sqcup\dots\sqcup S_{t}$, where $\Lambda_{i}=\Lambda_{j}$ (equivalently, $\boldsymbol{e}_{i}+\boldsymbol{e}_{j}\in{\rm Ker}\,\Lambda$) if and only if $i$ and $j$ lie in the same set $S_{k}$. Assume that there is a linear dependence $\Lambda_{j_{1}}+\dots+\Lambda_{j_{l}}=0$, where all the vectors $\{\Lambda_{j_{1}},\dots,\Lambda_{j_{l}}\}$ are pairwise different, that is $j_{p}\in S_{i_{p}}$, where $i_{p}\neq i_{q}$ if $j_{p}\neq j_{q}$. Then $\boldsymbol{e}_{j_{1}}+\dots+\boldsymbol{e}_{j_{l}}\in{\rm Ker}\,\Lambda$, and by (1)
@@
Indeed, $G_{1}\cap\dots\cap G_{M-1}$ is an (nM+1)(n-M+1)-RHDRHD. In particular, it is nonempty. But the intersection of any $n+1$ facets of $P$ is empty, hence $M-1\leqslant n$. Each $k$-face of $C(P,\Lambda)$ is a connected component of $G_{i_{1}}\cap\dots\cap G_{i_{n-k}}$, hence by Lemma 3.10 it is a $k$-RHDRHD if nk<Mn-k<M, and a $k$-RHSRHS if nk=Mn-k=M. ∎
@@
We will show that complexes $\mathcal{C}(P,\Lambda)$ and $\mathcal{C}(n,M)$ have the same combinatorial structure. Then an equivalence can be constructed inductively starting from the skeleton of minimal dimension, which is either the set of $(n+1)$ vertices for $M=n+1$, or (nM)(n-M)-sphere for $M<n+1$. Each face of greater dimension is a disk such that on its boundary a homeomorphism is already constructed. We extend the homeomorphism using the fact that the disk is the cone on its boundary.
For $n=2$ this is trivial.
@@
Let $n=3$. If $M=1$, then $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,1)$. If $M=2$, then $G_{1}$ and $G_{2}$ are $2$-disks such that $\partial G_{1}=\partial G_{2}=G_{1}\cap G_{2}$ is a circle. Hence, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,2)$. If $M=3$, then $G_{1}$, $G_{2}$ and $G_{3}$ are $2$-disks such that $G_{i}\cap G_{j}$ is a simple path for any $i\neq j$, and all these three paths have common ends, which form the 0-sphere $G_{1}\cap G_{2}\cap G_{3}$. Thus, the $1$-skeleton of $\mathcal{C}(P,\Lambda)$ is a theta-graph, and $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,3)$. If $M=4$, then each $G_{i}$ is a disk, the intersection $G_{i}\cap G_{j}$ of each two such disks is a simple path, the intersection of any three such disks is a point, and the intersection of all the four disks is empty. Then the $1$-skeleton of $\mathcal{C}(P,\Lambda)$ is a $K_{4}$-graph, and $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,3)$.
@@
Let $n=4$. If $M=1$, then $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,1)$. If $M=2$, then $G_{1}$ and $G_{2}$ are $3$-disks such that $\partial G_{1}=\partial G_{2}=G_{1}\cap G_{2}$ is a $2$-sphere. Hence, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,2)$. If $M=3$, then $G_{1}$, $G_{2}$ and $G_{3}$ are $3$-disks, such that $G_{i}\cap G_{j}$ is a $2$-disk for $i<j$, and $G_{1}\cap G_{2}\cap G_{3}$ is a circle. We have $G_{1}\cup G_{2}$ is a $3$-disk with a circle $G_{1}\cap G_{2}\cap G_{3}=\partial(G_{1}\cap G_{2})$ on its boundary, and the complement to this disk is the interior of $G_{3}$. Hence, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,3)$. If $M=4$, then $G_{1}$, $G_{2}$, $G_{3}$, $G_{4}$ are $3$-disks, $G_{i}\cap G_{j}$ is a $2$-disk for any $i<j$, $G_{i}\cap G_{j}\cap G_{k}$ is a simple path for any $i<j<k$, and all these four paths have the common ends, which form the 0-sphere $G_{1}\cap G_{2}\cap G_{3}\cap G_{4}$. Then $G_{1}\cup G_{2}\cup G_{3}$ is a $3$-disk with the theta-graph on the boundary formed by paths $G_{i}\cap G_{j}\cap G_{4}$, $i<j<4$. There is the path $G_{1}\cap G_{2}\cap G_{3}$ inside this disk connecting the vertices of the theta-graph and each $G_{i}\cap G_{j}$ is a $2$-disk bounded by the paths $G_{1}\cap G_{2}\cap G_{3}$ and $G_{i}\cap G_{j}\cap G_{4}$. The complement to $G_{1}\cup G_{2}\cup G_{3}$ in $\partial P$ is the interior of $G_{4}$. Thus, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,4)$. If $M=5$, then each $G_{i}$ is a $3$-disk, each intersection $G_{i}\cap G_{j}$, $i<j$, is a $2$-disk, each intersection $G_{i}\cap G_{j}\cap G_{k}$, $i<j<k$, is a simple path, each intersection $G_{i}\cap G_{j}\cap G_{k}\cap G_{l}$, $i<j<k<l$, is a point, and $G_{1}\cap G_{2}\cap G_{3}\cap G_{4}\cap G_{5}=\varnothing$. Then $G_{1}\cup G_{2}\cup G_{3}\cup G_{4}$ is a $3$-disk with a $K_{4}$-graph on its boundary formed by vertices $G_{i}\cap G_{j}\cap G_{k}\cap G_{5}$ and edges $G_{i}\cap G_{j}\cap G_{5}$. There is a point $G_{1}\cap G_{2}\cap G_{3}\cap G_{4}$ inside this disk such that the face structure in the disk $G_{1}\cup G_{2}\cup G_{3}\cup G_{4}$ is a cone over the face structure on its boundary with this apex. Also the complement to this disk is the interior of $G_{5}$. Thus, $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,5)\simeq\partial\Delta^{4}$.
@@
Let $\Lambda$ be a vector-coloring of rank $r$ of a simple polytope $P$ of dimension $n\leqslant 4$. Then $N(P,\Lambda)\simeq S^{n}$ if and only if $\mathcal{C}(P,\Lambda)\simeq\mathcal{C}(n,r)$.
@@ 4. Hyperelliptic manifolds $N(P,\Lambda)$
Let $\Lambda$ be a vector-coloring of rank $r$ of a simple $n$-polytope $P$, $n\leqslant 4$. Then an involution $\tau\in\mathbb{Z}_{2}^{r}$ is hyperelliptic (that is $N(P,\lambda)\simeq S^{n}$) if and only if $\tau$ is special (that is $\mathcal{C}(P,\Lambda_{\tau})\simeq\mathcal{C}(n,r-1)$, where $\Lambda_{\tau}$ is the composition of $\Lambda$ and the projection $\mathbb{Z}_{2}^{r}\to\mathbb{Z}_{2}^{r}/\langle\tau\rangle\simeq\mathbb{Z}_{2}^{r-1}$).
@@ 4.1. Hamiltonian subcomplexes
We call a subcomplex $\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ Hamiltonian, if each $q$-skeleton of $\mathcal{C}(P,c_{1})$ lies in the $(q+1)$-skeleton of $\mathcal{C}(P,c_{2})$.
@@
A subcomplex $\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ is Hamiltonian if and only if any two defining (n2)(n-2)-faces are disjoint.
@@
Let any two defining (n2)(n-2)-faces be disjoint. If $c_{2}(G_{i_{1}})=c_{2}(G_{i_{2}})=c_{2}(G_{i_{3}})$ and $i_{1}<i_{2}<i_{3}$, then either one of the intersections $G_{i_{1}}\cap G_{i_{2}}$ and $G_{i_{1}}\cap G_{i_{3}}$ is empty, or each intersection is a disjoint union of defining faces and any two faces from different unions are also disjoint. Then
@@
If $c_{2}(G_{i_{1}})=c_{2}(G_{i_{2}})\neq c_{2}(G_{i_{3}})=c_{2}(G_{i_{4}})$ and $i_{1}\neq i_{2}$, $i_{3}\neq i_{4}$, then either one of the intersections $G_{i_{1}}\cap G_{i_{2}}$ and $G_{i_{3}}\cap G_{i_{4}}$ is empty, or each intersection is a disjoint union of defining faces and any two faces from different unions are also disjoint. Then
@@
Now let the subcomplex be Hamiltonian. Consider two defining faces. If they are connected components of the same intersection $G_{i_{1}}\cap G_{i_{2}}$, then they are disjoint by definition. If they lie in different intersections $G_{i_{1}}\cap G_{i_{2}}$ and $G_{i_{3}}\cap G_{i_{4}}$, then there are two possibilities. First possibility is $c(G_{i_{1}})=c(G_{i_{2}})=c(G_{i_{3}})=c(G_{i_{4}})$. Then without loss of generality we can assume that $i_{3}\notin\{i_{1},i_{2}\}$. Hence, $G_{i_{1}}\cap G_{i_{2}}\cap G_{i_{3}}=\varnothing$ by Lemma 4.4(3) and $(G_{i_{1}}\cap G_{i_{2}})\cap(G_{i_{3}}\cap G_{i_{4}})=\varnothing$. The second possibility is $c(G_{i_{1}})=c(G_{i_{2}})\neq c(G_{i_{3}})=c(G_{i_{4}})$. Then $(G_{i_{1}}\cap G_{i_{2}})\cap(G_{i_{3}}\cap G_{i_{4}})=\varnothing$ by Lemma 4.4(3). This finishes the proof. ∎
@@
A set $\mathcal{S}=\{M_{1}^{n-2},\dots,M_{q}^{n-2}\}$ of pairwise disjoint (n2)(n-2)-faces of a complex $\mathcal{C}(P,c_{1})$ is a set of defining faces of some Hamiltonian subcomplex $\mathcal{C}(P,c_{2})\subset\mathcal{C}(P,c_{1})$ if and only if for each two facets $G_{i}\neq G_{j}$ of $\mathcal{C}(P,c_{1})$ admitting a sequence of facets $G_{i}=G_{i_{1}}$, $G_{i_{2}}$, $\dots$, $G_{i_{p}}=G_{j}$ such that for all $s$ the intersection $G_{i_{s}}\cap G_{i_{s+1}}$ contains a face from $\mathcal{S}$ either $G_{i}\cap G_{j}=\varnothing$ or each connected component of $G_{i}\cap G_{j}$ belongs to $\mathcal{S}$.